OpenStudy (anonymous):

How do you find average rate of change and identify a sequence based on a graph?

OpenStudy (misty1212):

HI! i guess it depends on the graph right?

OpenStudy (misty1212):

looks like you are dividing by 2 each time right?

OpenStudy (misty1212):

are you trying to find the sequence?

OpenStudy (anonymous):

yes, and i have to write the sequence out.

OpenStudy (anonymous):

and I have to find the average rate of change from n=0 to n=3

OpenStudy (misty1212):

2 gives 10 so 1 gives 20 and 0 gives 40

OpenStudy (misty1212):

sequence would be \[a_n=40\times \left(\frac{1}{2}\right)^n\]if you are starting at \(n=0\)

OpenStudy (misty1212):

average rate of change from 0 to 3 would be \(\frac{40-5}{3}\)

OpenStudy (anonymous):

great, thank you so much!